OpenStudy (anonymous):

The Sun burns about 4.4 × 1 0 6 tons of hydrogen per second. How much hydrogen does the Sun burn in one year? (Hint: one year = 3.16 × 1 0 7

OpenStudy (anonymous):

10 ^7

OpenStudy (anonymous):

10^6

OpenStudy (phi):

they want you to multiply two numbers that are in scientific notation. \[ 4.4 × 10^6 \frac{tons}{sec} \times 3.16 × 10^7 \frac{sec}{year} \]

OpenStudy (phi):

do you know how to do this?

OpenStudy (phi):

Here is an example by Khan http://www.khanacademy.org/video/scientific-notation-2?topic=worked-examples-4